Sally travels 990m south 876m west find the displacement and distance?
1
Expert's answer
2020-03-13T10:50:22-0400

The distance


"d=990+876=1866\:\rm m."

The displacement


"s=\sqrt{990^2+876^2}=1322\:\rm m."